OpenStudy (anonymous):

4(x-6)-(3x+2)=-5 how do i solve?

OpenStudy (anonymous):

distribute the 4 inside the prentesis then drop the prentesis then combine like terms

OpenStudy (anonymous):

so this would be 4x-24-3x+2=-5 x=17

OpenStudy (anonymous):

opp hold on

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

x=21

OpenStudy (anonymous):

isnt it 4x-3x then 24+2?

OpenStudy (anonymous):

it's 4x-24-3x-2=-5

OpenStudy (anonymous):

It's 4x-3x and then -24-2 that is x-26=-5 x=-5+26 x=21